OpenStudy (anonymous):

The length of a field is three times its width. If the perimeter is 64 m, what is the width of the field? A. 8 m B. 16 m C. 21 m D. 32 m

OpenStudy (anonymous):

the answer is c 21

OpenStudy (anonymous):

can i get a medal plz

OpenStudy (anonymous):

perimeter is all the sides added together l + l + w + w = 64 if the length (l) is three times the width (w) then l = w + w + w (or 3w) substitute that in for l above w+w+w + w+w+w + w + w = 64 add the w's , there are 8 so, 8w = 64 divide by 8 , w = 8 (64/8 = 8) the width is 8 the length is 3 * the width, or 3 * 8 = 24 check your work, 2 lengths, and 2 widths = 64 24 + 24 + 8 + 8 = 64 so the answer is A, 8m